# Finiteness criteria for the solutions of the sequence of semi-decomposable form equations and inequalities

### v1

@@ 3. Semi-decomposable form equations and inequalities
In order to prove Theorem 1.3, we first prove the following version of the Schmidt’s subspace theorem.
Theorem 3.1.
Let $k$ be a number field, $S$ a finite set of places of $k$ and $V$ an irreducible projective subvariety of ${\mathbb{P}}^{N}$ of dimension $m$ defined over $k$. Let $\{Q_{1},\ldots,Q_{q}\}$ be a family of $q$ homogeneous polynomials in $k[x_{0},\ldots,x_{N}]$ with the distributive constant $\Delta$ with respect to $V(\bar{k})$. Then for each $\epsilon>0$,
$\sum_{j=1}^{q}\sum_{v\in S}\dfrac{\lambda_{Q_{j},v}({\bf x})}{\deg Q_{j}}\leqslant\left(\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon\right)h({\bf x})$
for all ${\bf x}\in V(\bar{k})\setminus\bigcup_{i=1}^{q}\{Q_{i}=0\}$ but finitely many points.
We note that the inequality in Theorem 3.1 is weaker the one in Theorem 2.1. However, the conclusion in Theorem 3.1 is that such an inequality holds for all but finitely many points, while in Theorem 2.1 the conclusion is that the inequality holds only outside a proper Zariski-closed set.
Proof.
We see that $(\frac{m}{2}+1)^{2}\geqslant m+1$. Then, by applying Theorem 2.1 for the family $\{Q_{1},\ldots,Q_{q}\}$, we conclude that the set of ${\bf x}\in V(\bar{k})\setminus\bigcup_{i=1}^{q}\{Q_{i}=0\}$ with
(3.2) $\sum_{j=1}^{q}\sum_{v\in S}\log\left(\frac{\|{\bf x}\|_{v}^{\deg Q_{j}}\cdot\|Q_{j}\|_{v}}{\|Q_{j}({\bf x})\|_{v}}\right)^{1/{\deg Q_{j}}}>\left(\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon\right)h({\bf x})$
is contained in a finite union of proper subvariaties of $V(\bar{k})$. Note that these subvarieties are of dimension at most $m-1$.
Take $V^{\prime}$ be one of them with $\dim V^{\prime}\geqslant 1$. Let $\Delta^{\prime}$ be the distributive constant of the family $\{Q_{1},\ldots,Q_{q}\}$ with respect to $V^{\prime}$. Applying Theorem 2.1 to $V^{\prime}$ with the family $\{Q_{1},\ldots,Q_{q}\}$, we get that the set of ${\bf x}\in V^{\prime}\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ with
(3.3) $\sum_{j=1}^{q}\sum_{v\in S}\log\left(\frac{\|{\bf x}\|_{v}^{\deg Q_{j}}\cdot\|Q_{j}\|_{v}}{\|Q_{j}({\bf x})\|_{v}}\right)^{1/\deg Q_{j}}>(\Delta^{\prime}(\dim V^{\prime}+1)+\epsilon)h({\bf x})$
is contained in a finite union of proper subvarieties of $V^{\prime}$. Of course these subvarieties are of dimension at least $m-2$. Note that, from Remark 3.8 one has $\Delta^{\prime}\leqslant(m-\dim V^{\prime}+1)\Delta$ and hence
$\Delta^{\prime}(\dim V^{\prime}+1)\leqslant\Delta(m-\dim V^{\prime}+1)(\dim V^{\prime}+1)\leqslant\Delta\left(\frac{m}{2}+1\right)^{2}.$
Hence, if a point ${\bf x}\in V^{\prime}\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ satisfies (3.2) then it must satisfy (3.3). Therefore, the set of ${\bf x}\in V^{\prime}\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ satisfying (3.3) is contained in a finite union of proper subvariaties of $V(\bar{k})$ of dimension at least $m-2$. This argument holds for every such subvariety $V^{\prime}$. Hence, the set of ${\bf x}\in V(\bar{k})\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ satisfying (3.3) is contained in a finite union of proper subvariaties of $V(\bar{k})$ of dimension at least $m-2$.
By this way, after $m-1$ steps, we conclude that the set of ${\bf x}\in V(\bar{k})\setminus\bigcup_{i=1}^{q}\{Q_{i}=0\}$ satisfying (3.3) is contained in a finite union of proper subvariaties of $V(\bar{k})$ of dimension zero, i.e, this set is finite. The theorem is proved. ∎
@@
Suppose contrarily that there is an infinite sequence ${\bf x}_{n}=(x_{0,n},\ldots,x_{m,n})\in\mathcal{O}_{S}^{m+1}$ which satisfies (1.4) and (1.5). If the heights $\{h({\bf x}_{n})\}_{n\geqslant 1}$ are bounded then then the set $\{{\bf x}_{n};n\geqslant 1\}$ is actually a finite union of sets of proportional vectors. Therefore, we may suppose that $h({\bf x}_{n})$ are not bounded.
Without loss of generality, we may assume that $h({\bf x}_{n})\rightarrow\infty$ as $n\rightarrow\infty$. Then, by the assumption, (1.5) also holds. Furthermore it follows that $H_{S}({\bf x}_{n})\rightarrow\infty$ as $n\rightarrow\infty$ (since ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1}$). By $\max_{0\leqslant j\leqslant m}(Q_{j,n})\leqslant h(F_{n})+O(1)$, from the inequality (1.5) we deduce that
@@
Take $\epsilon$ a positive number small enough so that $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}+\lambda+\epsilon$. By Theorem 3.1, there is an infinite subsequence ${\bf x}_{n_{k}}\in\mathcal{O}_{S}^{m+1},k=1,2,\ldots$, of $\{{\bf x}_{n}\}$, which without loss of generality we may assume to be $\{{\bf x}_{n}\}$ itself, such that
@@
$\prod_{v\in S}\frac{\|{\bf x}_{n}\|_{v}^{l}\cdot\prod_{j=1}^{q}\|Q_{j,n}\|_{v}}{\|F_{n}({\bf x}_{n})\|_{v}}\leq(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}$
@@
$\frac{H_{S}^{l}({\bf x}_{n})\cdot\prod_{v\in S}\prod_{j=1}^{q}\|Q_{j,n}\|_{v}}{\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}}\leq(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}.$
@@
$H_{S}^{l}({\bf x}_{n})$ $\leqslant\frac{1}{c^{\prime}}\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}\cdot(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}$
@@
a) In [8], Q. Ji and Q. Yan and G. Yu consider the case where each semi-decomposable form $F_{n}$ can be factorized into moving hypersurfaces in general position. Therefore, once we can define the notion of “distributive constant” for families of moving hypersurface with respect to a subvariety of ${\mathbb{P}}^{m}$ defined over a number field $k$ then we may generalize the result of Ji, Yan and Yu.
b) With the same arguments in [8] (see also [3]), we will get a result on semi-decomposable form equation as follows.
Consider the $S$-integer solutions of a sequence of semi-decomposable form equations of the forms
@@
Applying Theorem 1.3 for $c$ and $l-d\Delta\left(\frac{m}{2}+1\right)^{2}-1<\lambda<l-d\Delta(\frac{m}{2}+1)^{2}$, we get the following theorem.
@@
Let $\ell,m$ be positive integers. Let $k^{\prime}$ be a finite extension of $k$ and $S\subset$ $M_{k}$ a finite set containing all archimedean places. For $n=1,2,\ldots$, let $F_{n}({\bf x})=$ $F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ denote a sequence of semi-decomposable forms of degree $\ell$ with the distributive constant not exceeding a positive number $\Delta$ for each $n$. Assume that $F_{n}=Q_{1,n}\cdots Q_{q,n}$ over $k^{\prime}$ with $\deg Q_{j,n}=d_{j},1\leqslant j\leqslant q$ and $\Delta_{\{Q_{j,n}\}_{1\leq j\leq q}}\leq\Delta$ for each $n$. Let $d=\max_{1\leqslant j\leqslant q}d_{j}$. Assume that $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}$. Let $G_{n}({\bf x})\in\mathcal{O}_{S}[{\bf x}]$ such that $\deg G_{n}<\ell-d\Delta\left(\frac{m}{2}+1\right)^{2}$ for each $n$. Then there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-non-proportional ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1}$ for which
@@
and $\text{and }h(F_{n})$ $=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
In particular, for each nonzero $S$-integer $b$, the equation
$F(x_{0},\ldots,x_{m})=b.$
(under the given assumption in Corollary 1.6 for $F$) has finitely many $\mathcal{O}_{S}^{*}$-nonproportional solutions, provided $\deg F>d\Delta\left(\frac{m}{2}+1\right)^{2}$ with $d=\max_{1\leqslant j\leqslant q}d_{j}$.

### v2

@@ 4. Semi-decomposable form equations and inequalities
Suppose on the contrary that there is an infinite sequence ${\bf x}_{n}=(x_{0,n},\ldots,x_{m,n})\in\mathcal{O}_{S}^{m+1}$ which satisfies (1.2) and (1.3). If the heights $\{h({\bf x}_{n})\}_{n\geqslant 1}$ are bounded then the set $\{{\bf x}_{n};n\geqslant 1\}$ is actually a finite union of sets of proportional vectors. Therefore, we may suppose that $h({\bf x}_{n})$ are not bounded.
Without loss of generality, we may assume that $h({\bf x}_{n})\rightarrow\infty$ as $n\rightarrow\infty$. Then, by the assumption, (1.3) also holds. Furthermore it follows that $H_{S}({\bf x}_{n})\rightarrow\infty$ as $n\rightarrow\infty$ (since ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1}$). By $\max_{0\leqslant j\leqslant m}h(Q_{j,n})\leqslant h(F_{n})+O(1)$, from the inequality (1.3) we deduce that
@@
Take $\epsilon$ a positive number small enough so that $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}+\lambda+\epsilon$. By Theorem 3.4, there is an infinite subsequence ${\bf x}_{n_{p}}\in\mathcal{O}_{S}^{m+1},p=1,2,\ldots$, of $\{{\bf x}_{n}\}$, which without loss of generality we may assume to be $\{{\bf x}_{n}\}$ itself, such that
@@
$\prod_{v\in S}\frac{\|{\bf x}_{n}\|_{v}^{\ell}\cdot\prod_{j=1}^{q}\|Q_{j,n}\|_{v}}{\|F_{n}({\bf x}_{n})\|_{v}}\leq(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}$
@@
$\frac{H_{S}^{\ell}({\bf x}_{n})\cdot\prod_{v\in S}\prod_{j=1}^{q}\|Q_{j,n}\|_{v}}{\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}}\leq(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}.$
@@
$H_{S}^{\ell}({\bf x}_{n})$ $\leqslant\frac{1}{c^{\prime}}\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}\cdot(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}$
@@
a) If the family of moving hypersurfaces is in general position, our result will imply [9, Theorem 1.2] of Ji, Yan and Yu.
b) With the same arguments in [9] (see also in [4]), we will get a result on semi-$q$-decomposable form equations as follows.
Consider the $S$-integer solutions of a sequence of semi-decomposable form equations of the form
@@
Applying Theorem 1.1 for $c$ and $\ell-d\Delta\left(\frac{m}{2}+1\right)^{2}-1<\lambda<\ell-d\Delta(\frac{m}{2}+1)^{2}$, we get the following theorem.
@@
Let $\ell,m,q$ be positive integers. Let $k^{\prime}$ be a finite extension of $k$ and $S\subset$ $M_{k}$ a finite set containing all archimedean places. For $n=1,2,\ldots$, let $F_{n}({\bf x})=$ $F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ denote a sequence of semi-$q$-decomposable forms of degree $\ell$. Assume that $F_{n}=Q_{1,n}\cdots Q_{q,n}$ over $k^{\prime}$ with $\deg Q_{j,n}=d_{j},1\leqslant j\leqslant q$ and $\{Q_{1,n},\ldots,Q_{q,n}\}$ has the distributive constant not exceeding a positive number $\Delta$ for each $n$. Let $d=\max_{1\leqslant j\leqslant q}d_{j}$. Assume that $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}$. Let $G_{n}({\bf x})\in\mathcal{O}_{S}[{\bf x}]$ such that $\deg G_{n}<\ell-d\Delta\left(\frac{m}{2}+1\right)^{2}$ for each $n$. Then there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-non-proportional ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1}$ for which
@@
$\text{and }h(F_{n})$ $=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
In particular, for each nonzero $S$-integer $b$, the equation $F(x_{0},\ldots,x_{m})=b$ (under the given assumption in Corollary 1.4 for $F$) has finitely many $\mathcal{O}_{S}^{*}$-nonproportional solutions, provided $\deg F>d\Delta\left(\frac{m}{2}+1\right)^{2}$ with $d=\max_{1\leqslant j\leqslant q}d_{j}$.